On March 1, 2021, Gold Examiner receives $160,000 from a local bank and promises to deliver 94 units of certified 1-oz. gold bars on a future date. The contract states that ownership passes to the bank when Gold Examiner delivers the products to Brink’s, a third-party carrier. In addition, Gold Examiner has agreed to provide a replacement shipment at no additional cost if the product is lost in transit. The stand-alone price of a gold bar is $1,410 per unit, and Gold Examiner estimates the stand-alone price of the replacement insurance service to be $90 per unit. Brink’s picked up the gold bars from Gold Examiner on March 30, and delivery to the bank occurred on April 1. Required: 1. How many performance obligations are in this contract? 2. to 4. Prepare the journal entry Gold Examiner would record on March 1, March 30 and April 1.
Answer #1
1
Number of performance obligations in the contract is 2.
2
Debit Credit
March 01, 2021 Cash 160000
     Deferred revenue—gold bars 150400
     Deferred revenue—insurance 9600
March 30, 2021 Deferred revenue—gold bars 150400
        Sales revenue 150400
April 01, 2021 Deferred revenue—insurance 9600
       Service revenue 9600
Workings:
Value of the gold bars 132540 =1410*94
Stand­alone selling price of the insurance 8460 =90*94
Total of stand­alone prices 141000
Allocation:
Deferred revenue—gold bars 150400 =160000*132540/141000
Deferred revenue—insurance 9600 =160000*8460/141000